Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 654     Entry time: 06 Oct 2025, 09:45
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 06.10.2025 
General Announcement:
- M. Doering announces on the LIMP-25: 
1. Pressurizer exchange is postponed from today to a date likely in two weeks, to be informed; 
2. Magnet cabling in XTD10 starts on Wednesday and pre-absorber work must finish before (see VAC/XRO); 
3. News on timing system (see EEE).

CTRL:
- Preparing for Karabo3 rollout as scheduled.

DA:
- Nothing to report.

EEE:
- Update of timing system is scheduled for today after lunch. Groups are informed.
- PLC side: SPB and MID tests are ongoing.

VAC:
- XTD10 pre-absorber works: XPD part is done and VAC has to extract the pre-absorber and seal the vacuum, no issues to do in the next two days to avoid collision with magnet cabling.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Nothing to report.

XRO:
- Nothing to report; pre-absorber work is ongoing well.

LAS:
- Nothing to report; no collisions with timing update.

UNSYS:
- Nothing to report.

SPB:
- Nothing to report.

FXE:
- Nothing to report.

MID:
- All fine at MID.

HED:
- Nothing to report.

SCS:
- Nothing to report.

SQS:
- Nothing to report.

SXP:
- All fine at SXP.
ELOG V3.1.4-7c3fd00